# Class Unbiasing for Generalization in Medical Diagnosis

### v1

@@ 1 Introduction
Medical data sets are often biased, which compromises the accuracy and reliability of models trained using empirical risk minimization (ERM). A common bias is class imbalance, where models tend to prioritize the majority class at the expense of the minority classes, leading to poor performance on the latter. To mitigate this issue, researchers have proposed various methods, such as oversampling or undersampling classes [1, 2], or using class weights to adjust the loss function [1, 3, 4]. Apart from the class imbalance mentioned in Section 1, there are other invisible biases that arise from spurious features, including patient-specific characteristics (e.g., age [7], gender [7, 8], or speaker identity [3]) and environment-specific factors (e.g., equipment types [7], equipment manufacturers [7], or variations of diagnosis equipment across hospitals [9]). To address these issues, researchers recommend using balanced and diverse training sets and reducing the influence of spurious features in model predictions [8, 3, 7].
@@
Notably, the existence of class-feature bias is a major issue in medical diagnosis, where the primary procedure lies in identifying the presence or absence of specific biomarkers or their measured values that distinguish between disease and non-disease states. Therefore, leveraging class-shared features that are informative to both disease and non-disease states can help increase diagnosis accuracy and understand disease mechanisms. Conversely, neglecting these shared features risks overlooking critical indicators of disease progression or response to treatment, potentially leading to misdiagnosis, delayed treatment, or inadequate monitoring. In addition, class-feature bias can be even more problematic under class imbalance scenarios, where one class dominates the dataset, further exacerbating the issue of biased performance. Therefore, developing models that can effectively learn and leverage class-shared features is essential for improving the accuracy and effectiveness of medical diagnoses.
@@
Leveraging insights from information theory, the key intuition to mitigate the class-feature bias is to ensure the features learned by the model are informative for both classes. Specifically, the target-related information in the features for each class can be measured by class-wise cross-entropy classification losses ($\mathcal{L}^{c}$ and $\mathcal{L}^{\neg c}$). Therefore, a class loss inequality ($\mathcal{L}^{c}\neq\mathcal{L}^{\neg c}$) indicates that the model relies on features that are non-informative to both classes, which can be a sign of existence of class-feature bias under class-balance. Meanwhile, class loss inequality can be also caused by class-imbalance, where the model may under-utilize information from features for the minority class, leading to lower class-wise classification loss. Motivated by [24], we propose a class-wise inequality loss $\mathcal{L}_{\text{cls-ineq}}$ that equalizes the classification losses contributed by the samples from the positive and negative classes, respectively. The class-wise inequality loss can help learn a class-unbiased model (Cls-unbias) that mitigates both class imbalance and class-feature bias simultaneously. To improve the effectiveness of the class-wise inequality loss under class imbalance, we combine the cross-entropy loss with a class-wise group distributionally robust optimization (G-DRO) objective [5], which adaptively upweights the high-loss class.
To the authors’ best knowledge, this study is the first to identify and treat the class-feature bias as a research problem in deep-learning based medical diagnosis, empirically demonstrating its detrimental effect on binary classification performance. We provide both theoretical and empirical evidence showing the effectiveness of the class-wise inequality loss in addressing class-feature biases in balanced and imbalanced datasets. Extensive experiments were conducted on synthetic and real-world datasets. Our results demonstrate that promoting class-unbiased classification consistently improves model performance across diverse tasks and datasets with different input modalities.
2 Results
2.1 Toy Experiment: Demonstrating and Mitigating the Class-feature Bias
We designed simple experiments using linear classifiers to demonstrate class-feature bias in binary classification under both class-balanced and class-imbalanced conditions. In this section, we illustrated an example of class-specific and class-shared features, and showed that both class imbalance and class-feature bias can negatively impact classification performance of typical ERM (trained using cross-entropy loss only), and that making the classifier unbiased can improve generalization to out-of-distribution scenarios.
Consider a 2D space containing feature vectors $\bm{x}$’s, where $\bm{x}=[f_{1},f_{2}]^{\top}$ is an input feature vector containing two features $f_{1}$ and $f_{2}$, and $y\in\{0,1\}$ is the classification label of $\bm{x}$. We set $f_{1}\in\{a,b\}$ and $f_{2}\in\{a,b\}$, where $a\sim\mathcal{N}(0,1)$ and $b\sim\mathcal{N}(3,1)$. Therefore, we may express $\bm{x}=[(a|b),(a|b)]^{\top}$, where $(a|b)$ means selecting either $a$ or $b$ from the set $\{a,b\}$. We applied this principle to generate three datasets, namely $D^{\text{tr}}$, $D^{\text{test1}}$, and $D^{\text{test2}}$. We used $D^{\text{tr}}$ to train a linear classifier and tested it on $D^{\text{test1}}$ and $D^{\text{test2}}$. Fig. 1 and Fig. 2 show the probability distributions of features $f_{1}$ and $f_{2}$ for the positive class ($y=1$) and negative class ($y=0$) under class-balanced and class-imbalanced scenarios, respectively.
Class-shared Feature. In Fig. 1, $f_{1}$ is equally effective in classifying $\bm{x}$’s into the positive or negative classes for all datasets ($\mathcal{D}^{\text{tr}}$, $\mathcal{D}^{\text{test1}}$, and $\mathcal{D}^{\text{test2}}$), meaning that it is invariant to the differences in three datasets. Also, $f_{1}$ is an example of class-shared features mentioned in Section 1, because it can differentiate both the positive and negative classes.
Class-specific Feature. Fig. 1(a) shows that $f_{2}$ is a class-specific feature in $\mathcal{D}^{\text{tr}}$. More specifically, for a classifier that uses $f_{2}$ as the only feature, an unknown sample having $f_{2}\ll 3$ (say $f_{2}=-2$) indicates that it is likely a sample from the positive class (red). Meanwhile, assigning samples with $f_{2}\approx 3$ to the negative class (blue) can ensure that all negative samples are correctly classified. Therefore, $f_{2}$ is specific to the negative class. On the other hand, Fig. 1(b) and Fig. 1(c) show that $f_{2}$ is an irrelevant feature in $\mathcal{D}^{\text{test1}}$ and $\mathcal{D}^{\text{test2}}$, as its value does not determine the class of any unknown samples in these datasets.
Spurious Feature. Applying the model trained in $\mathcal{D}^{\text{tr}}$ to the data in $\mathcal{D}^{\text{test2}}$ reveals that $f_{2}$ is a spurious feature. This is because although $f_{2}$ can partly determine the class labels in the samples in $\mathcal{D}^{\text{tr}}$, it fails to do the same in $\mathcal{D}^{\text{test2}}$.
In the ideal class-unbiased condition, the classofier should learn a vertical decision boundary in Fig. 1 and Fig. 2 to separate the two classes. This means that the classifier pays full attention to $f_{1}$, while assigning no attention to $f_{2}$, resulting in high accuracy on both $\mathcal{D}^{\text{test1}}$ and $\mathcal{D}^{\text{test2}}$. However, the actual decision boundaries of the linear classifier trained by the ERM (Eq. 1) and the proposed class-unbiased method (Eq. 9) on balanced and imbalanced datasets are shown in Fig. 1 and Fig. 2, respectively. This discrepancy between the ideal and actual decision boundaries suggests the existence of class-specific features and spurious features.
Table 1 shows the classification performance of the linear classifier on the three datasets shown in Fig. 1 and Fig. 2 under class-balanced and class-imbalanced scenarios, respectively. We report the accuracy for the positive and negative classes separately (referred to as pos-acc and neg-acc, respectively) to assess class-wise performance. The classifier comprises two input nodes and one output node, with the normalized weights of the two input features indicated in the last column. [^1] We omit the bias term of the classifier as it does not affect our discussion. To simulate the class-imbalanced and class-balanced situations, we set $p(y=1)=0.1$ in the class-imbalanced experiments (Rows 3 and 4) and $p(y=1)=0.5$ in the class-balanced experiments (Rows 1 and 2) in the training set. Meanwhile, we kept $p(y=1)=p(y=0)=0.5$ in the test sets for all experiments.
The key conclusions from the toy experiment are summarized below:
@@ 2.1 Toy Experiment: Demonstrating and Mitigating the Class-feature Bias
    The classifier trained by ERM (Row 1) relies heavily on the class-specific feature ($f_{2}$), as evidenced by the relatively large weights assigned to it. This reliance leads to degraded performance on $\mathcal{D}^{\text{test1}}$ for the positive class. The performance degradation becomes more severe under the class-imbalanced condition, as evident in Row 3. Worse, as mentioned earlier, because $f_{2}$ is a spurious feature for $\mathcal{D}^{\text{test2}}$, its performance on the biased class (negative class) can be significantly compromised (Row 1).
@@
    By comparing the weights in Row 3 with those in Row 1, we observe that class imbalance increases the model’s reliance on the unintended feature $f_{2}$. This leads to improved detection of the negative samples but significantly degrades the performance on the positive samples. The model in Row 3 exhibits the combined effects of class-specific bias and class imbalance, which jointly distort the learned decision boundaries and amplify the asymmetric performance between the two classes.
@@
    In all situations, Cls-unbias, which promotes the reduction of class loss inequality, can effectively alleviate the reliance on the class-specific feature $f_{2}$, while simultaneously improving performance. Moreover, it helps mitigate the class imbalance problem.
(a) $\mathcal{D}^{\text{tr}}$
(b) $\mathcal{D}^{\text{test1}}$
(c) $\mathcal{D}^{\text{test2}}$
Figure 1: The decision boundaries of a linear classifier trained by the ERM (Eq. 10) and the proposed class-unbiased method (Eq. 9) under the class-balance scenario. The contourplots show distributions of the two features ($f_{1}$ and $f_{2}$) for the positive (pos) and negative class (neg) for the three datasets: $\mathcal{D}^{\text{tr}}$, $\mathcal{D}^{\text{test1}}$, and $\mathcal{D}^{\text{test2}}$. The positive and negative classes have equal prior in $\mathcal{D}^{\text{tr}}$.
(a) $\mathcal{D}^{\text{tr}}$
(b) $\mathcal{D}^{\text{test1}}$
(c) $\mathcal{D}^{\text{test2}}$
Figure 2: The decision boundaries of a linear classifier trained by the ERM (Eq. 10) and the proposed class-unbiased method (Eq. 9) under the class-imbalance scenario. The contourplots show distributions of the two features ($f_{1}$ and $f_{2}$) for the positive (pos) and negative class (neg) for the three datasets: $\mathcal{D}^{\text{tr}}$, $\mathcal{D}^{\text{test1}}$, and $\mathcal{D}^{\text{test2}}$. The prior probability of the positive class is 0.1 in $\mathcal{D}^{\text{tr}}$, i.e., $P(y=1)=0.1$.
Table 1: Results for the experiments on the synthetic example. For the imbalanced dataset, $P(y=1)=0.1$. For meaningful comparisons of the attention to the features $f_{1}$ and $f_{2}$ by different models, we normalized the model’s weights $\bm{w}=[w_{1},w_{2}]$ by $\bm{w}/\|\bm{w}\|$ and show them in the “Normalized Weights” column.
Row Method Class-balanced Train Test 1 Test 2 Normalized Weights
neg-acc pos-acc neg-acc pos-acc neg-acc pos-acc
1 ERM 0.848 0.850 0.880 0.703 0.679 0.825 [0.858,
0.514]
2 Cls-unbias 0.853 0.864 0.876 0.780 0.802 0.845 [0.962,
0.275]
3 ERM 1.000 0.001 1.000 0.000 1.000 0.004 [0.663,
0.749]
4 Cls-unbias 0.852 0.850 0.880 0.708 0.770 0.841 [0.925,
0.379]
Although simplified, this toy experiment gives an example of class-shared and class-specific features, clearly demonstrates both the existence of the class-feature bias and the effectiveness of our method in addressing it. These findings motivate the application of our approach to more complex, real-world scenarios, as discussed in subsequent sections.
2.2 Class Unbiasing in Real-world Datasets
To evaluate the real-world effectiveness of our class-unbiased method, we tested it on five high-dimensional datasets across speech and medical imaging: two speech-based datasets for depression detection (DAIC-WOZ [26] and MODMA [29]), one for speech-based dementia diagnosis (ADReSS [31]), and two imaging datasets [32]—BreastMNIST for breast cancer detection and a binarized version of RetinaMNIST for diabetic retinopathy detection using retinal fundus images. Performance is reported using the macro-averaged F1-score (MF1), which provides a reliable measure under class-imbalanced conditions. The proposed method is compared against two baseline approaches: standard Empirical Risk Minimization (ERM) and class-weighted ERM (ERM (cls-w)).
The main results across five datasets are presented in Table 2. Overall, the proposed class unbiased model consistently outperforms the baselines, across all datasets.
2.3 Ablation Study
In medical datasets, class imbalance often results in unreliable modeling of data distributions. The results on the DAIC-WOZ dataset (Table 3) demonstrate a case of severe class imbalance, where the model trained with ERM fails to learn meaningful patterns. Specifically, the model collapses into always predicting the majority (negative) class, thereby effectively minimizing the cross-entropy loss and achieving around 70% accuracy on the training set. Apparently, this seemingly decent performance is an illusion, as the model fails to detect the minority (positive) class. While the variants of ERM partially alleviate this issue—either through a class-weighting strategy (ERM (cls-w)) or by placing equal emphasis on all classes (ERM (per-cls))—the overall performance remains poor, indicating that the model is still negatively affected by class imbalance and potential class-feature bias.
Class-wise group distributionally robust optimization ($\mathcal{L}_{\text{g-dro}}$) can be used to alleviate the effect of class imbalance, as it adaptively emphasizes the under-explored, underperforming class by assigning higher weight to the minority class. As shown in Table 3, solely using class-wise group distributionally robust optimization (Row 3) is a better approach than using the variants of ERM (ERM (per-cls) & ERM (cls-w)) to handle the class-imbalance problem in the DAIC-WOZ dataset, as the former promotes the exploration of under-represented classes to reduce class loss inequality. However, it fails to handle class-feature bias, because its strategy is to maximize exposure to the minority class rather than minimize reliance on biased class-specific features. In contrast, the class-unbiased models (Rows 4 & 5) substantially improve performance by reducing class loss inequality, thereby mitigating both class imbalance and class-feature bias through explicitly promoting class-wise loss equality, which leads to more generalizable and reliable performance.
Despite the proposed class-inequality loss can reduce class-imbalance, its effectiveness can be negatively affected by severe class-imbalance. This is because the unreliable modeling of class-wise information, caused by class imbalance, can make equalizing class-wise classification losses misleading. Therefore, under scenarios where class imbalance significantly impacts training, we recommend using $\mathcal{L}_{\text{g-dro}}$ to obtain a better initial estimate of class-wise information, providing a stronger reference for optimizing the class-inequality loss ($\mathcal{L}_{\text{cls-ineq}}$).
Comparing Row 4 and Row 5 of Table 3 further reveals additional improvement brought by incorporating $\mathcal{L}_{\text{g-dro}}$, demonstrating its effectiveness in helping the class-wise inequality loss achieve its intended effect. This is particularly evident on the DAIC-WOZ dataset, where severe class imbalance skews the model’s estimation of class-wise information, thereby limiting the effectiveness of $\mathcal{L}_{\text{cls-ineq}}$. By introducing $\mathcal{L}_{\text{g-dro}}$, we mitigate the distortion in the estimated class-wise information caused by class imbalance, improving the overall performance.
DAIC-WOZ MODMA ADReSS BreastMNIST RetinaMNIST
[…]
Table 3: Ablation study on DAIC-WOZ.
Row Method Valid Test
pos-F1 neg-F1 MF1 Acc pos-F1 neg-F1 MF1 Acc
[…]
2.4 Case Study on Balanced and Imbalanced Datasets.
(a) $\mathcal{L_{\text{cls-ineq}}}$
(b) $\mathcal{L}^{\text{pos}}$ and $\mathcal{L}^{\text{neg}}$.
(c) Classification loss $\mathcal{L}_{\text{ERM}}$ on validation set.
(d) MF1 on validation set.
(a) $\mathcal{L_{\text{cls-ineq}}}$
(b) $\mathcal{L}^{\text{pos}}$ and $\mathcal{L}^{\text{neg}}$.
(c) Classification loss $\mathcal{L}_{\text{ERM}}$ on validation set.
(d) MF1 on validation set.
To empirically investigates the effect of the proposed class-unbiased approach on balanced and imbalanced datasets, we resampled the RetinaMNIST dataset to construct two settings: a balanced dataset with 100 samples (50% positive) and an imbalanced dataset with 100 samples (25% positive). The experimental results are shown in Table 4. The training processes on both balanced and imbalanced datasets are shown in Figures 3 and 4, respectively.
Row Methods cls-balanced MF1 Acc
[…]
3 Discussion
3.1 From Binary Classification to Multi-class Classification
Extending Cls-unbias from binary classification to multi-class classification is feasible. Similar to [24], one possible strategy is to reduce the variance of the class-wise losses. Encouraging class-wise loss equality remains effective when the number of classes is small. Determining the number of classes for which the multi-class extension of Cls-unbias remains valid can be left to future work.
Finally, we emphasize that Cls-unbias remains highly useful and important in medical diagnosis settings, i.e., binary or few-class classification, where strict feature selection and extraction are critical, as classification failures can lead to severe consequences, including misdiagnosis, delayed treatment, or inadequate monitoring.
3.2 Ethical Impact caused by Class-feature Bias
Class-feature bias, while often subtle and invisible without deliberate awareness, poses significant ethical concerns, especially as deep learning models are increasingly adopted in sensitive domains such as medical diagnosis. This bias arises when the models’ potential reliance on features that are strongly correlated with only a subset of classes, leading to biased performance and poor generalization on other classes. Critically, these “other classes” are more likely to be misclassified, leading not only to increased costs in time and healthcare resources due to misdiagnosis but also to potential ethical harms. For instance, as discussed in the Section 1, using BMI as a dominant predictive feature for diabetes may lead the model to incorrectly associate high BMI with the disease, reinforcing harmful stereotypes and potentially marginalizing individuals with atypical clinical presentations. Such biased representations can exacerbate healthcare disparities and erode trust in AI-assisted medical decisions. Therefore, it is essential to identify and mitigate class-feature bias during model development. Furthermore, this calls for a broader integration of causal learning approaches to ensure more stable, fair, and ethically responsible generalization across diverse patient populations.
4 Method
4.1 Definition of Class-feature Bias
@@ 4.3 Group Distributionally Robust Optimization for Class-imbalance
4.4 Baselines: Empirical Risk Minimization
where $\mathcal{B}$ denotes a batch of data and $\ell(\bm{x},y)$ is the loss (e.g., cross-entropy loss for classification) for the input-target pair $(\bm{x},y)$.
Variants of ERM, such as class-weighted ERM and per-class ERM, help mitigate the impact of the dominant classes under class-imbalance [6]. In particular, class-weighted ERM modifies the loss function by applying class-dependent scaling factors $\alpha_{y}$’s to adjust for class imbalance:
4.5 Related Work on Domain Generalization
The proposed method relates to domain generalization, which aims to improve model robustness under distributional shifts—where the distribution of unseen test data differs from that of the training data. Existing approaches to address the distributional shift mainly fall into two complementary categories: 1) reduce domain discrepancies by making the data representations from different domains similar, using techniques such as data augmentation [10, 11, 12], feature alignment [13, 14, 15], or gradient matching [16, 17]; and 2) remove domain-specific information/factors, typically through disentanglement [18, 19], invariant risk minimization [20, 21, 22], parameter decomposition [23], or sample reweighting [5]. For example, UndoBias [23] enables support vector machines (SVMs) to handle dataset bias by learning both shared and dataset-specific weights, allowing the SVMs to separate dataset-specific biases from general object features. Invariant risk minimization (IRM) [20, 22] aims to learn predictors that generalize across diverse environments by promoting invariance in the learned representations.
@@ 4.6.1 Speech Datasets
DAIC-WOZ. We utilized the Distress Analysis Interview Corpus Wizard-of-Oz (DAIC-WOZ) dataset [26], a widely-used benchmark for depression diagnosis. The dataset comprises clinical interviews from 189 participants, totaling approximately 30 hours of conversation data. Each interview was transcribed and annotated with the participants’ depression status, as determined by the PHQ-8 scale [27]. According to the official division of DAIC-WOZ dataset, we trained the models on the training set, validated on the development set, and tested on the test set. Pretrained wav2vec features [28] were extratced and preprocessed as stated in [3].
MODMA. The Multimodal Open Dataset for Mental Disorder Analysis (MODMA) [29] is a depression dataset collected and released by Lanzhou University from the Second Hospital of Lanzhou University, containing audio recordings from 52 individuals. In MODMA, the positive class constitutes 43% of the data. We followed the splits of training, validation, and test set in [30]. The waveforms were segmented into 3.84-second segments, and 80-dimensional filterbank features (FBanks) were extracted as the model input.
ADReSS. The ADReSS dataset [31] is a speech-based dementia dataset, comprising speech recordings and their transcripts of participants describing the Cookie Theft picture from the Boston Diagnostic Aphasia Exam. The dataset includes 78 dementia participants and 78 healthy controls. Audio recordings were segmented to 3.84s segments, and 80-dimensional FBanks were extracted for training.

### v2

@@ 1 Introduction
Medical data sets are often biased, which compromises the accuracy and reliability of models. A common bias is class imbalance, where models tend to prioritize the majority class at the expense of the minority classes, leading to poor performance on the latter. To mitigate this issue, researchers have proposed various methods, such as oversampling or undersampling classes [1, 2], or using class weights to adjust the loss function [3, 1, 4].
@@
Notably, the existence of class-feature bias is a major issue in medical diagnosis, where the primary procedure lies in identifying the presence or absence of specific biomarkers or their measured values that distinguish between disease and non-disease states. Leveraging class-shared features that are informative to both disease and non-disease states can help increase diagnosis accuracy and understand disease mechanisms. Conversely, neglecting these shared features risks overlooking critical indicators of disease progression or response to treatment, potentially leading to misdiagnosis, delayed treatment, or inadequate monitoring. Therefore, developing models that can effectively learn and leverage class-shared features is essential for improving the accuracy and effectiveness of medical diagnoses.
@@
Class-feature bias can be even more problematic under class imbalance scenarios, where one class dominates the dataset, further exacerbating the issue of biased performance. In this study, we aim to train a class-unbiased model (Cls-unbias) that mitigates both class imbalance and class-feature bias simultaneously. Leveraging insights from information theory, the key intuition is to ensure the features learned by the model are informative for both classes. Specifically, we propose a class-wise inequality loss that equalizes the classification losses contributed by the samples from the positive and negative classes, respectively. To improve the effectiveness of the class-wise inequality loss under class imbalance, we combine the cross-entropy loss with a class-wise group distributionally robust optimization (G-DRO) objective [5], which adaptively upweights the high-loss class.
We summarize our contributions as follows:
@@
    To the authors’ best knowledge, this study is the first to treat the class-feature bias as a research problem in medical diagnosis, empirically demonstrating its detrimental effect on binary classification performance.
@@
    We provide both theoretical and empirical evidence showing the effectiveness of the class-wise inequality loss in addressing class-feature biases in balanced and imbalanced datasets.
@@
    For practical considerations, we propose to use G-DRO to enhance the inequality loss’s performance under class imbalance.
  4. 4.
    We conducted extensive experiments on synthetic and real-world datasets. Our results demonstrate that promoting class-unbiased classification consistently improves model performance across diverse tasks and datasets with different input modalities.
2 Background & Related Work
2.1 Empirical Risk Minimization
where $\mathcal{B}$ denotes a batch of data and $\ell(\bm{x},y)$ is the loss (e.g., cross-entropy loss for classification) for the input-target pair $(\bm{x},y)$. Variants of ERM, such as class-weighted ERM and per-class ERM, help mitigate the impact of the dominant classes under class-imbalance [6]. In particular, class-weighted ERM modifies the loss function by applying class-dependent scaling factors $\alpha_{y}$’s to adjust for class imbalance:
2.2 Bias in Medical Diagnosis
Inherent biases in medical datasets are often invisible, causing models trained by ERM to produce inaccurate predictions. Apart from the class imbalance mentioned in Section 1, there are other biases that arise from spurious features, including patient-specific characteristics (e.g., age [7], gender [7, 8], or speaker identity [3]) and environment-specific factors (e.g., equipment types [7], equipment manufacturers [7], or variations of diagnosis equipment across hospitals [9]). To address these issues, researchers recommend using balanced and diverse training sets and reducing the influence of spurious features in model predictions [8, 3, 7]. However, class-feature bias has largely been overlooked. This paper highlights the problem of class-feature bias in medical diagnosis and proposes a solution.
2.3 Domain Generalization
Domain generalization aims to improve model robustness under distributional shifts, where the distribution of unseen test data differs from that of the training data. Existing approaches to address the distributional shift mainly fall into two complementary categories: 1) reduce domain discrepancies by making the data representations from different domains similar, using techniques such as data augmentation [10, 11, 12], feature alignment [13, 14, 15], or gradient matching [16, 17]; and 2) remove domain-specific information/factors, typically through disentanglement [18, 19], invariant risk minimization [20, 21, 22], parameter decomposition [23], or sample reweighting [5]. For example, UndoBias [23] enables support vector machines (SVMs) to handle dataset bias by learning both shared and dataset-specific weights, allowing the SVMs to separate dataset-specific biases from general object features. Invariant risk minimization (IRM) [20, 22] aims to learn predictors that generalize across diverse environments by promoting invariance in the learned representations.
3 Class-feature Bias
@@ 4.2 Group Distributionally Robust Optimization for Class-imbalance
5 Experimental Setup
We evaluated the effectiveness of the class unbiased method on high-dimensional data (speech and image) in the real world. We provide a comprehensive analysis on the impact of class unbiasing on the class imbalance problem and the class-feature bias problem.
@@ 5.1.1 Speech Datasets
DAIC-WOZ
We utilized the Distress Analysis Interview Corpus Wizard-of-Oz (DAIC-WOZ) dataset [26], a widely-used benchmark for depression diagnosis. The dataset comprises clinical interviews from 189 participants, totaling approximately 30 hours of conversation data. Each interview was transcribed and annotated with the participants’ depression status, as determined by the PHQ-8 scale [27]. According to the official division of DAIC-WOZ dataset, we trained the models on the training set, validated on the development set, and tested on the test set. Pretrained wav2vec features [28] were extratced and preprocessed as stated in [3].
MODMA
The Multimodal Open Dataset for Mental Disorder Analysis (MODMA) [29] is a depression dataset collected and released by Lanzhou University from the Second Hospital of Lanzhou University, containing audio recordings from 52 individuals. [^1] https://modma.lzu.edu.cn/data/index/ In MODMA, the positive class constitutes 43% of the data. We followed the splits of training, validation, and test set in [30]. The waveforms were segmented into 3.84-second segments, and 80-dimensional filterbank features (FBanks) were extracted as the model input.
ADReSS
The ADReSS dataset [31] is a speech-based dementia dataset, comprising speech recordings and their transcripts of participants describing the Cookie Theft picture from the Boston Diagnostic Aphasia Exam. The dataset includes 78 dementia participants and 78 healthy controls. Audio recordings were segmented to 3.84s segments, and 80-dimensional FBanks were extracted for training.